a)solve dy/dx - 3ye^3x =y. Given that y=1 when x=0



b)find the solution of sinx dy/dx + y cosx = sin^3 x cosx


1
Expert's answer
2022-02-07T13:59:55-0500

a)


"y'-3ye^{3x}=y"

"y'=y(1+3e^{3x})"

"\dfrac{dy}{y}=(1+3e^{3x})dx"

Integrate


"\int \dfrac{dy}{y}=\int (1+3e^{3x})dx"

"\ln (|y|)=x+e^{3x}+\ln C"

"y=Ce^{x+e^{3x}}"

"1(0)=1"


"1=Ce^{0+e^{3(0)}}=>C=e^{-1}"

"y=e^{x+e^{3x}-1}"

b)


"d(y\sin x)=\sin^3x\cos xdx"

Integrate


"\int d(y\sin x)=\int \sin^3x\cos xdx"

"y\sin x=\dfrac{\sin^4 x}{4}+C"

"y=\dfrac{\sin^3 x}{4}+\dfrac{C}{\sin x}"